When determining the energy of a reaction where 0.250 mol of a substance with a Hcomb = -2250.0 kJ/mol reacts, we perform this calculation:


-2250.0 kJ/1 mol = x kJ/0.250 mol OR 0.250 mol (-2250.0 kJ/ 1 mol)


Why can't we simply say that the energy of the reaction is -2250.0 kJ?


1
Expert's answer
2021-09-03T01:46:59-0400

The first way is more appropriate. The energy released when 1 mol of the substance reacts = - 2250.0 KJ So, the energy released when 'x' mol of the substance. You should calculate the enthalpy per 0.250 mol, so you have to calculate ΔHcomb × n = -2250 × 0.250 = -562.5 (J).